OpenStudy (anonymous):

simplify (w/6)^2 a)w/36 b)w^2/6 c)w^2/36 d)36w^2

OpenStudy (anonymous):

\[\left( \frac{ w }{ 6 } \right)^2 \implies \frac{ w^2 }{ 6^2 }\] when you square something, and it's in brackets as such, it implies everything in the expression is squared. :) Hope that makes sense! Now try finishing it off.

OpenStudy (mertsj):

\[(\frac{w}{6})^2=\frac{(w)^2}{(6)^2}=\frac{w^2}{36}\]

OpenStudy (anonymous):

simplify: (2m^3n^2)^5 a)10m^8n^7 b)10m^15n^10 c)32m^8n^7 d)32m^15n^10

OpenStudy (zale101):

Show your attempt on this one

OpenStudy (anonymous):

i think C

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Can you show your work, on how you got that?

OpenStudy (anonymous):

i did 2*2*2*2*2=32 5+3=8 5+2=7

OpenStudy (anonymous):

Remember you have to multiple the variables as well.